OpenStudy (anonymous):

Helpful steps please, A coil of wire rotating in a magnetic field induces a voltage given by e = 20 sin ((piT/4)-(pi/2)). where t is time in seconds. Find the smallest positive time to produce a voltage of 10 square root 3. ( would this be 3.33pi sec?)

OpenStudy (perl):

did you solve 20 sin ((piT/4)-(pi/2)) = 10 sqrt(3)

OpenStudy (anonymous):

I have sqrt(3)/2

OpenStudy (perl):

looks good so far sin( pi /4 * t - pi/2 ) = sqrt(3)/2 we know that when theta = Pi/3, sin (Pi/3) = sqrt(3)/2 so solve Pi/3 = Pi/4 * t - Pi/2 Pi/2 + Pi/3 = Pi/4 * t

OpenStudy (anonymous):

oh so its just 3.33sec?